OpenStudy (anonymous):

The area of an isosceles trapezoid is 90 in squared, the height is 6 inches, and the congruent legs are 10 inches. Find the length of the longer base.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

take out the area that they give you and your left with a triangle right?

OpenStudy (amistre64):

90 - 60 =30 we have an area of 30 to account for

OpenStudy (anonymous):

I dont think im following you. How'd you get 60?

OpenStudy (amistre64):

the height is 6 so....... 6b/2 = 30 3b = 30 b = 10 more than the other one

OpenStudy (amistre64):

the area of a rectangle is base times height right? 6.10 = 60

OpenStudy (amistre64):

90 total - 60 that is given = 30 left

OpenStudy (anonymous):

But 10 is the leg not a base. So it would make triangle not a square right?

OpenStudy (amistre64):

the longer base = 20

OpenStudy (amistre64):

i see it lol

OpenStudy (anonymous):

Im confused. What is 90 and how did you know that 10 was the base>>

OpenStudy (amistre64):

then we pythag it to find the 'base'

OpenStudy (anonymous):

see what?

OpenStudy (amistre64):

i see that i read the problem wrong and know how to fix it...

OpenStudy (amistre64):

6^2 + b^2 = 10^2 b^2 = 100 - 36 b = sqrt(64) = 8 , roght?

OpenStudy (amistre64):

like this; and the problem says the trap area = 90

OpenStudy (anonymous):

Hold on im working it out

OpenStudy (anonymous):

Right i got that

OpenStudy (amistre64):

6*16 ----- = 48 for the 'area' of the triangle 2 90 - 48 = 42 6(7) = 42...... small side = 7 large side = 23 then right?

OpenStudy (anonymous):

im working it

OpenStudy (anonymous):

Wait I think I'm getting it but your saying 48 is the area for both the small triangles right?

OpenStudy (amistre64):

yes;

OpenStudy (amistre64):

together they make one triangle with a height of 6 and a base of 16

OpenStudy (anonymous):

Oh! I got it. THanks for the help